A sphere has a mass of 150 kg and radii of 0.4 m. Assuming it starts from rest, how much more work is required to get this sphere rolling on a horizontal surface at 2.5 m/s than to get it sliding at 2.5 m/s on a frictionless horizontal surface.

Please show all work and include explanations! Thanks!
Answer #1

When rolling, we need to consider linear and rotational kinetic energies.

When sliding, we need to consider only linear kinetic energy.

NOTE - You did not mention if the sphere is solid or hollow.

I am assuming it to be a solid sphere.

-------------------------------------------------------------------------------------------

While rolling,

W = 1/2mv2 + 1/2Iw2

w = v/r

I = 2/5 * m * r2 ( For solid sphere)

W = 468.75 + 1/2 * 2/5 * 150 * 0.42 * (2.5/0.4)2

W = 468.75 + 187.5

W = 656.25 J

---------------------

Now, when its only sliding

W = 1/2 mv2

W = 468.75

Therefore,

we need 187.5 J of more work to get this sphere rolling on a horizontal surface at 2.5 m/s than to get it sliding at 2.5 m/s on a frictionless horizontal surface.